Base your answer to que records of the Dickinson in inventory: stion 9 on the following information taken from the Office Supply Company, pertaining to calculators it had Total Price $80 81 81 81 84 86 Amount Beginning inventory Jan. 1, 20x3 Purchase Jan. 8, 20X3 Purchase Jan. 13, 20X3 Purchase Jan. 18, 20X3 Purchase Jan. 22, 20X3 Units 40 30 30 40 30 30 $3,200 2,430 2,430 3,240 2,520 2,580 Purchase Jan. 27, 20X3 9. There were 30 calculators on hand in inventory as of January 31, 20X3. Which of the following is the correct inventory valuation, using the average cost method A. $2,400 B. $2,460 C. $2,580 D. $2,610
Answer #1

9) The correct answer is B because...

The average cost method takes the average price of all the items

Purchase date (20X3) Number of units price per unit($) Amount($)
Jan 1( begining inventory) 40 80 3200
jan 8 (purchase) 30 81 2430
jan 13 (purchase) 30 81 2430
jan 18 (purchase) 40 81 3240
jan 22 (purchase) 30 84 2520
jan 27 (purchase) 30 86 2580

The total amount = $16400

Total Number of units = 200

Therefore the average price = 16400/200 = $82 per unit

Given that the inventory left is 30 units, so the value of the inventory is = 30*82 = $2460

Thank you..